### v1

Abstract
Fifteen years of the Fermi Large Area Telescope (LAT) data in the halo region of the Milky Way (MW) are analyzed to search for gamma rays from dark matter annihilation. Gamma-ray maps within the region of interest (, ) are modeled using point sources, the GALPROP models of cosmic-ray interactions, isotropic background, and templates of Loop I and the Fermi bubbles, and then the presence of a halo-like component is further examined. A statistically significant halo-like excess is found with a sharp peak around 20 GeV, while its flux is consistent with zero below 2 GeV and above 200 GeV. Examination of the fit residual maps indicates that a spherically symmetric halo component fits the map data well. The radial profile agrees with annihilation by the smooth NFW density profile, and may be slightly shallower than this, especially in the central region. Various systematic uncertainties are investigated, but the 20 GeV peak remains significant. In particular, the halo excess with a similar spectrum is detected even relative to the LAT standard background model, which does not depend on GALPROP or other model templates. The halo excess can be fitted by the annihilation spectrum with a mass 0.5–0.8 TeV and annihilation cross section (5–8) for the channel. This cross section is larger than the upper limits from dwarf galaxies and the canonical thermal relic value, but considering various uncertainties, especially the density profile of the MW halo, the dark matter interpretation of the 20 GeV “Fermi halo” remains feasible. The prospects for verification through future observations are briefly discussed.
1 Introduction
Even though nearly a century has passed since Fritz Zwicky inferred the existence of dark matter from the motion of galaxies in the Coma cluster [1], dark matter remains one of the greatest mysteries in astrophysics, cosmology, and fundamental physics [2, 3, 4]. Weakly interacting massive particles (WIMPs) in the mass range of 1 GeV to 100 TeV have long been considered one of the leading candidates, because they interact only through the weak interaction, and because the thermal relic amount with the annihilation cross section expected from particle physics is close to the observed dark matter density [3, 4, 5, 6]. The existence of WIMPs can be probed with direct searches by underground experiments, indirect searches using annihilation products (e.g., gamma-rays and neutrinos), and by particle accelerator experiments.
Regarding indirect searches by gamma-rays, the sensitivities of the Fermi Large Area Telescope (Fermi LAT) and ground-based air Cherenkov gamma-ray telescopes have reached a level where the intensity expected from the canonical annihilation cross section can be detected. When WIMPs directly annihilate into two photons, the line gamma-ray emission is easy to distinguish, but the branching ratio into line emission is highly uncertain. In the more mainstream annihilation mode, the Standard Model particles such as or are first generated, followed by a cascade that produces gamma rays with a continuum energy spectrum. The spectrum still shows a broad peak around a specific energy related to the WIMP mass, which contrasts with power-law spectra of many astrophysical objects. Numerous explorations have been conducted toward various astronomical objects or regions in search of such signals, including the Milky Way (MW) Galactic center (GC) [7, 8, 9], dwarf galaxies [10, 11, 12, 13] and dark subhalos [14, 15] in the MW halo, the Andromeda galaxy [16, 17], galaxy groups and clusters [18, 19, 20, 21], and the extragalactic background radiation [22, 23, 24, 25].
This study will search for annihilation gamma rays from the MW halo, avoiding the GC and disk regions. Although the annihilation radiation spreading across the MW halo is not as strong as that near the GC, it is possible to avoid the influence of gamma rays originating from various objects and cosmic rays around the GC and the disk, thereby potentially increasing the effective signal-to-noise ratio. Looking back on past searches towards the MW halo [26, 27, 28, 29, 30, 31, 32], it seems that interest in this region for a dark matter search has waned compared to other objects or regions since the early 2010s. This is probably because the so-called GC GeV excess with a sharp morphological peak at the GC, which may be a result of dark matter annihilation, has been found around a photon energy of a few GeV [33, 34, 7, 8]. Another possible reason is that halo-scale (up to 50 deg from the GC) diffuse radiation structures, such as the Fermi bubbles [35, 36, 37], have been found. They are unlikely to originate from dark matter and are confusing when searching for an annihilation signal in the MW halo region.
However, the GC GeV excess can also be explained by astrophysical phenomena rather than dark matter, such as millisecond pulsars [8]. Then the true dark matter signal may appear in other photon energy ranges. The dark matter interpretation of the GC excess suggests that the radial density profile near the GC is with 1.1-1.3, which is more centrally concentrated than the standard Navarro-Frenk-White (NFW) [38] profile ( in the central region). If the true density profile is the NFW or less centrally concentrated, dark matter signals may be more easily detected in the halo region than at the GC. In 2012, the Fermi LAT team analyzed the MW halo region in search of dark matter using two years of the LAT data, which is much smaller than the amount accumulated to date. From that time until now, understanding and analysis methods for diffuse radiation in the halo region, such as the Fermi bubbles and Loop I, have also matured. For these reasons, it would be worthwhile to investigate the presence or absence of a halo-like radiation in addition to the known components based on the latest LAT data. This study reports the results of a dark matter search using 15 years of the LAT data.
The structure of this paper is as follows. Section 2 describes the LAT data used in this work, the region of interest in the analysis, the likelihood analysis method, and various model map templates considered. The results of the search for a halo-like excess are presented in section 3. Section 4 discusses the interpretation of the results in terms of dark matter annihilation, and section 5 presents an overview and some discussion.
2 Analysis methods
2.1 The LAT data and the region of interest
This work is based on the Pass 8 UltraClean class of the 15-year public Fermi LAT data from 2008 August 4 to 2023 August 2. In addition, a zenith angle cut of is adopted to reduce the influence of radiation from the Earth’s atmosphere. To avoid deterioration of angular resolution in the low photon energy range, this study analyzes only the range above 1 GeV. The angular resolution (68% containment angles) of the LAT is 0.85∘, 0.16∘, and 0.10∘ at 1, 10, and 100 GeV, respectively111https://www.slac.stanford.edu/exp/glast/groups/canda/lat_Performance.htm [39]. The data are divided into 13 energy bins, equally spaced in logarithmic space, with the minimum and maximum energy bin centers at 1.51 and 814 GeV, respectively.
The region of interest (ROI) in this study is set to and around the GC, and the disk area of is excluded from the dark matter search. Photon count maps and exposure maps with pixel scales of in Cartesian coordinates are created from the public Fermi LAT data and tools222https://fermi.gsfc.nasa.gov/ssc/data/, resulting in 960960 pixels in the ROI including the disk region.
2.2 Likelihood analyses
In many studies fitting a model map to the LAT data, the photon counts and the model expectations are compared in each pixel. However, the halo-like component is expected to be much smoother than the pixel scale of 0.125∘. Therefore, we divide the ROI into cells with a width of and calculate the likelihood based on photon counts and expectations in these cells. This not only speeds up calculations, but also prevents the likelihood from being unnecessarily biased by pixel-scale fine discrepancies between observed counts and model predictions. The maximum likelihood fitting is performed independently for each photon energy bin. Let be the observed photon count in the -th cell and be the number of photons expected from the model, and then the likelihood function based on Poisson statistics is
where the sum is for all cells. The expected count is calculated by summing up the expected counts over all pixels in the cell:
where represents each pixel, is the total model diffuse photon flux per unit photon energy and solid angle to the angular direction of the -th pixel, and are the LAT exposure (in units of ) and the solid angle of the -th pixel, and is the width of the photon energy bin under consideration. Though the expected counts should be convolved with the point spread function (PSF), this process is omitted, since the cell size is sufficiently larger than the LAT angular resolution in the photon energy range considered here.
The model diffuse flux is a linear combination of various map templates:
where is the template map of the -th model component. Then the best-fit values and statistical errors of the fitting parameters are obtained by maximizing the likelihood function using the Markov chain Monte Carlo (MCMC) method.
2.3 Model map templates
Following the analyses on the Fermi bubbles and the GC GeV excess by the Fermi LAT team [37, 7], the following components are considered to fit the gamma-ray maps in the ROI: (1) point sources, (2) diffuse interstellar emission by cosmic-ray interactions, (3) isotropic radiation, (4) Loop I, (5) the Fermi bubbles, and (6) a halo-like emission. The modeling of these components will be explained in order below. However, since the Fermi bubble template will be created during the fitting process in this study, it will be described in section 3.1.
The template of point sources in the LAT 14-year Source Catalog333https://fermi.gsfc.nasa.gov/ssc/data/access/lat/14yr_catalog/ (4FGL-DR4,
gll_psc_v35.fit) [40, 41] is calculated using the parametrized spectra from the catalog and the energy-dependent LAT point spread function (PSF).
4FGL-DR4 also provides a catalog of extended sources, in addition to the point sources.
Since the extended sources are not many, circular regions with a radius twice that of the semimajor axis of the catalog are masked and excluded from the analysis.
Gamma rays produced by cosmic rays propagating in the MW are the most significant component of the Galactic diffuse gamma-ray emission.
Here, we utilize the GALPROP code444https://galprop.stanford.edu/webrun/ (ver. 54) [42, 43, 44], which numerically solves the generation and propagation of cosmic rays and calculates the gamma-ray intensity resulting from interactions with interstellar gas and radiation fields.
Among the various GALPROP parameter sets used in the LAT team’s study of the Galactic diffuse emission [45], SLZ6R30T150C2 is adopted as the baseline model in the present work.
This parameter set is used for the inverse Compton scattering (ICS) component in the Galactic interstellar emission model (GIEM, gll_iem_v07.fits)555https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html [46, 47] recommended by the LAT team for normal point source analysis.
The GALPROP model radiation is divided into two components, gas and ICS, which are fitted independently.
The gas component is the sum of pion-decay gamma rays produced by cosmic-ray hadrons, and bremsstrahlung from cosmic-ray electrons and positrons.
The GALPROP model emission can be further divided into more components, and the fitting results with such a larger number of sub-components or different model parameters will be discussed in section 3.4.3.
Loop I is a giant structure spanning more than 100∘ in the radio sky [48], and although it is unclear whether it is nearby supernova remnants or an outflow from the GC [49, 50], it has also been detected by LAT [51]. We adopt the geometric model of Loop I, which assumes a uniform emissivity within two shells with radii of 50–100 pc, based on the nearby hypothesis. The parameters of the geometric model are taken from the LAT team papers [37, 7], which differ slightly from the values in Wollben’s original paper [49]. The normalization factors for each of the two shells are treated as independent fit parameters ().
To calculate the halo component model map, we first assume that the density profile of the MW dark matter halo is NFW, and use the parameters from the Via Lactea II simulation [52, 53]. Therefore, the density profile within the virial radius = 402 kpc is , where , 21 kpc, and . The distance from the GC to the solar system is set at 8 kpc. Annihilation gamma-ray diffuse flux is proportional to the line-of-sight integration of the gamma-ray emissivity per unit volume. If dark matter is distributed smoothly, the emissivity obeys , and the map template in this case is denoted as NFW-. According to numerical simulations, fine substructures and subhalos exist within halos composed of cold dark matter, and subhalos with high internal densities may dominate annihilation gamma-ray luminosity. If the number density of subhalos is proportional to the smooth density profile, the gamma-ray emissivity would be . This emissivity also applies when dark matter particles decay rather than annihilate. Therefore, the map template of (NFW-) will also be tested. Subhalos may be destroyed by strong tidal forces near the center of the parent halo [52], in which case emissivity in the central region would be suppressed, but this is not tested in this study. Finally, based on previous studies indicating that the density profile of () is preferred for the GC GeV excess, the model template of (NFW-), which becomes the same emissivity profile as the GC excess in the central region using the original NFW, is included in the fit for this component.
The map morphologies of the point sources and GALPROP have photon energy dependencies determined by the models, while other model maps do not depend on photon energy. The initial values of in the MCMC chain are the same as the original normalization for the models of point sources and GALPROP. The initial of the isotropic component is set to , which is typical of the reported fluxes of the extragalactic gamma-ray background radiation [54]. For other components, the initial values of the MCMC chain are set to zero. During the MCMC fitting, the fitting parameters are limited to the range for known components such as point sources, the GALPROP models, the isotropic component, and Loop I, but is allowed for the halo components, which are explored as unknown components.
3 Search for a halo-like excess
3.1 Construction of the Fermi bubble template
Unlike Loop I and the MW halo components, there are no parameterized analytical models for the Fermi bubbles. Therefore, the map template for the bubbles used in this study is made here, mostly following the methods adopted in the LAT team’s papers [37, 7]. As the first step, a flat (i.e., isotropic) template within the rough area of the Fermi bubbles is assumed, and it is fit to the data together with the other model components. Here, the GC GeV excess is included in the fit as a known component, but other halo components are not considered. At this stage, the fit is performed in the whole ROI, including the Galactic disk region (), to ensure an accurate fit to the GC excess. Looking at the best-fit image of the bubbles (i.e., sum of the flat template and fit residuals, smoothed by the Gaussian with ), the energy-independent boundary of the flat template is iteratively improved.
The image of the Fermi bubbles obtained in this way is shown in figure 1, together with the final boundary of the flat template, in the two energy bins of 1.5 and 4.3 GeV, which are in the typical photon energy range where the Fermi bubbles are prominent. Since the template of the Fermi bubbles will only be used at in subsequent halo component searches, the Galactic disk region () is not shown. Residuals of bright point sources are visible in some places, and these are likely due to uncertainties in the catalog spectral models and PSF. The fit in this study is performed on a cell-by-cell basis with the cell size of , and PSF errors should have little effect on the results.
The best-fit spectra (calculated by the best-fit in each energy bin) for all the model components by this fit are shown in figure 2. Here, it should be noted that the spectral flux shown in this figure excludes the disk region for comparison with the later analyses, though the model fitting is performed including the disk region. The spectra of the model templates are generally consistent with the results of similar analyses by the LAT team [7], indicating that there are no serious problems or errors in this analysis. In particular, the peak photon energy and flux of the GC GeV excess ( at 1∘ from the GC at 2.5 GeV) are also consistent with those reported in the literature [8].
The morphology of the Fermi bubbles is also consistent with what has been reported in the literature. However, in the northern region ( 10–15∘), negative residuals are clearly visible, especially at the lower energy bin of 1.5 GeV. This feature corresponds to the regions of particularly strong flux in the gas component map of GALPROP (see Figure 13 in section 3.3 later), and is likely due to discrepancies between the GALPROP model and the data. A similar feature is seen in the LAT team’s analysis using GALPROP (Fig. 22 of ref. [37]), and the LAT team used the method of spectral component analysis to produce a bubble template without this feature. Since the purpose of the present study is not to create an accurate template of the Fermi bubbles, the following approach will be taken to explore an additional halo component.
Among the two energy bins, the 4.3 GeV map is chosen as the source to create the template, because the Fermi bubbles are more clearly seen. In the map shown in Figure 1, the positive flux regions are dominated by the Fermi bubbles. Therefore, the map of positive regions, regardless of whether it is inside or outside the boundaries of the flat template, will be used as the energy-independent template for the bubbles. On the other hand, in order to eliminate the effects of discrepancies between GALPROP and the data, the map of the negative flux regions is introduced into the fit as another template. Since these two probably have different energy spectra, they are included as two independent templates.
With this treatment, the model and data maps should agree well without additional model components at around 4.3 GeV. Even if a halo component is present near this energy bin, it will be absorbed by these templates. However, we are interested in a halo-like emission with a characteristic photon-energy dependence. If such a component really exists, we expect to see positive or negative excesses of the model halo template in other energy regions. This is the signal we will be searching for.
3.2 Fitting results with a halo-like component
From here, the fits will be performed excluding the Galactic disk region () to search for a halo-like component. The flat template of the Fermi bubbles is replaced with the positive and negative residual templates obtained in the previous section. The fit parameter is limited to for the positive residual template (the Fermi bubbles), but not for the negative template, because the inverted negative template (i.e., positive flux) may be favored for a better fit, if the true morphology of the Fermi bubbles is relatively flat and the discrepancy between GALPROP and the data decreases in higher energy bins,
First, the fit is performed without the GC excess or any other halo component, and the best-fit spectra are shown in figure 3. The spectrum of the new Fermi bubble template (positive residual map) is almost the same as that of the flat template, and as expected, the negative template spectrum is softer than the Fermi bubbles and closely resembles that of the gas component. Another notable point is that the ICS component is softer than the fit including the Galactic plane (figure 2). This suggests that, compared to the GALPROP model, the ICS spectrum softens from the Galactic plane to high Galactic latitudes.
Next, the fit results with the GC GeV excess component (NFW-) added are shown in figure 4. The GC excess flux is the same as the previous fit that includes the Galactic plane (figure 2) at 4.3 GeV. This is how it should be, because the two residual templates are created at this energy bin. However, in higher photon energy regions, the GC excess component rises sharply and peaks at around 20 GeV, where the flux is more than 10 times higher than the previous fit. This means that extrapolating this fit to the GC would significantly exceed the observed data. These results indicate that a halo-like component exists around 20 GeV, with a density profile shallower than that assumed for the GC GeV excess, and is prominent in the region.
Therefore, to test a flatter density profile, we perform fits using the NFW- and NFW- halo models instead of the GC excess, and the results are shown in figures 5 and 6, respectively. Only one of the three halo models is used in a fit to avoid degeneracy between the similar halo maps. Since the best-fit halo flux becomes negative at the lowest photon energy bin, figure 7 shows the spectra of the three halo models on a linear plot. The flatter the density profile, the higher the mean background flux in the ROI, but in all cases, the intensity of the halo components rises sharply with photon energy between 1 GeV and 20 GeV, reaching a statistically significant (13–19) peak at the 21 GeV bin. At higher energy bins, the halo flux decreases rapidly, becoming almost zero at 200 GeV or higher.
The halo components are not completely zero at the 4.3 GeV bin, where the positive and negative templates are created from the residual map, because the GC excess component was included in the fit that yielded the residual map. The halo component flux decreases toward lower photon energies, becoming negative at 1.5 GeV. This may be attributed to the difference between the GALPROP gas component and the data, which becomes more pronounced at lower energies. Another possibility is that part of the halo component at 4.3 GeV has been absorbed in the residual map templates. If the halo component weakens toward lower energy below 4.3 GeV, the fit of the halo component could be negative to cancel out the halo component absorbed in the residual map templates.
Figure 8 illustrates the likelihood (relative to the no-halo fit) of the fits with three different halo models, to compare the goodness of fit. In the 2.5 GeV bin, the halo components are fitted to be nearly zero (figure 7), and a deep dip is seen in the figure. Around the peak photon energy of the halo excess (10–20 GeV), the flatter NFW- model fits better than NFW- at 95% confidence level (CL). On the other hand, the shallowest NFW- model fits even worse than the NFW- model (at a significance much higher than 95% CL), suggesting that a profile close to NFW- best fits the data in the region of .
3.3 Morphology and radial profile of the halo-like excess
Figure 9 shows the radial angular profiles of three halo model fits, at the energy bin of 21 GeV where the halo component is the most prominent. The diffuse flux of the best-fit halo model is shown as a curve, and the data points are the sum of the best-fit model and the fit residuals, in a ring region centered on the GC within the ROI (, ). This plot confirms what we found in the previous section, namely that the NFW- model best fits the data in terms of likelihood. The density profile of the NFW- model is steeper than the data, while the NFW- model is shallower. For comparison, the best fit flux of the GC GeV excess (the NFW- model in the fit including the Galactic plane) is also shown in the figure as the curve of “NFW- ()”. The halo component must have a shallow profile that does not exceed this curve when extrapolated to the GC. Though the NFW- model best fits at , it is several times higher than the GC excess at around 1∘ from the GC. This suggests that the angular profile that best fits the data across the entire ROI is slightly flatter than NFW-, or that the angular profile becomes gradually flatter than NFW- toward the GC.
To verify how well the halo model matches the data in terms of morphology, map images are shown in figure 10, for the NFW- fit at the 21 GeV bin. When the fit is made without a halo model component, the fit tries to minimize the halo component, and hence, the fit would be biased if the halo component physically exists. However, a halo-like residual can be seen even in such a fit (top left). In the fit including NFW- as the halo model, the map of the model plus residual (i.e., the observed map with all model maps except for the halo subtracted) shows a clearer halo-like structure (top right), which is spherically symmetric and consistent with the model map (bottom left). The fit residual with the halo model (bottom right) is significantly reduced compared to the fit without the halo (top left), indicating that the halo fit is successful overall in the ROI.
To examine the fits in energy bins other than the peak (21 GeV), figures 11 and 12 show maps of the NFW- model plus residual at bins lower and higher than 21 GeV, respectively. These maps show the trend seen in figure 7, with the negative fit at the lowest 1.5 GeV, nearly zero at 2.5 GeV, the halo component becoming stronger toward 21 GeV, and then rapidly weaker beyond 21 GeV. The morphology of the halo model plus residual is generally spherically symmetric and does not appear to contradict the profile of the assumed model halo.
Finally, the images of the best-fit model map templates considered in the fits, other than the halo, are shown in figure 13 for comparison. The morphology of the GALPROP gas component and Loop I differ greatly from the halo model, and it is unlikely that they are degenerate with each other. The ICS components of GALPROP are shown separately for each of the three target photon fields: optical (direct radiation from stars), infrared (emission from interstellar dust grains), and the cosmic microwave background (CMB) radiation. These are shown separately here to check morphology, but it should be noted that they are treated as a single component in the fitting of the baseline analysis. Though the optical and infrared ICS components show a morphology most similar to that of the halo, they are considerably flattened along the Galactic plane compared with the halo. The possible degeneracy between the ICS and halo components will be discussed in section 4.1.
3.4 Examination of systematic uncertainties
Here, we examine systematic uncertainties concerning the evidence for the halo-like excess on the following points: (1) uncertainties in the Fermi bubble template, (2) treatment of point sources, (3) model parameters and number of free components in GALPROP, and (4) halo-like excess search using the Fermi LAT standard background model (GIEM) as the reference.
3.4.1 The Fermi bubble template
The Fermi bubble template was created from the positive residual map at 4.3 GeV; however, a portion of the halo component in that energy bin may be absorbed by the template, resulting in an underestimate of the halo component in the fit. To check the dependence on the choice of energy bin in this process, an alternative fit result is shown in Figure 14 (top), in which the positive (the Fermi bubbles) and negative residual templates are replaced with those of the 1.5 GeV map shown in figure 1. As expected, the energy bin where the halo component becomes zero has shifted to the lower side, and the halo flux at the 20 GeV peak has increased, but there is no change in the basic shape of the spectrum with a peak near 20 GeV.
As shown in figure 1, due to the strong negative residual around 10–15∘, the positive residual template may be biased from the true structure of the Fermi bubbles. To test the dependence on the Fermi bubble template structure, the fit result obtained by replacing the positive residual template with the simple flat template is shown in Figure 14 (top). The halo component has decreased compared to the baseline analysis, and this may be because the baseline analysis attempted to compensate for the concavity caused by the negative residual with the halo component. However, the overall spectral shape, including the rising rate of the halo component from 1 to 20 GeV, remains unchanged. The excess at the 21 GeV bin is still statistically significant (9.1).
3.4.2 Point sources
In the baseline analysis, point sources are subtracted using the spectral models in the 4FGL-DR4 catalog and the LAT PSF. Instead, the fit result after masking circular areas around the point sources with a radius equal to the 95% containment angle is shown in figure 14 (top). The result is almost unchanged from the baseline analysis. Due to the deterioration of the LAT angular resolution, the area to be masked increases toward lower photon energies, resulting in a large statistical error at the lowest bin.
3.4.3 The GALPROP models
In the study of the Galactic diffuse emission, the LAT team used 128 sets of various combinations of the five GALPROP parameters [45]. In the present study, five cases are tested in which the value of one parameter among the five is changed from the parameter set used in the baseline analysis. The baseline GALPROP parameter set SLZ6R30T150C2 means the Lorimer [55] distribution of pulsars for the cosmic-ray source distribution, the vertical and radial boundaries of the cosmic-ray confinement region and kpc, the spin temperature K for the opacity of the 21 cm line, and the magnitude cut at 2 mag for the reddening by dust as a gas tracer. Figure 14 (bottom) shows the results of five different fits in which these parameters are changed as follows: (1) the supernova remnant (SNR) distribution [56] for cosmic-ray sources, (2) 10 kpc, (3) 20 kpc, (4) K, and (5) the magnitude cut 5 mag for . In most cases, the fit results are almost unchanged from the baseline analysis, but relatively large changes are observed when the cosmic-ray source distribution is changed to SNR. In this case, the halo component is positively fitted at all energy bins and is nearly zero at the lowest energy bin. This suggests that one reason for the negative halo fits at the lowest energies in the baseline analysis may be the uncertainty of the gas component of GALPROP.
In the LAT team paper on the GC GeV excess [7], the GALPROP gas component is divided into five regions of Galactocentric radius (0–1.5, 1.5–3.5, 3.5–8, 8–10, and 10–50 kpc), and the ICS component is divided into the three sub-components of the different target photon radiation fields (optical, infrared, and CMB), which are fitted as independent templates. To check the dependence on the number of free GALPROP components, the NFW- fit is performed with the same treatment about GALPROP as ref. [7], and the result is shown in Figure 14 (top). The halo component decreases in bins below 12 GeV from the baseline analysis, but remains almost unchanged at the 21 GeV bin, and the overall trend of the spectrum peaking at 20 GeV remains unchanged.
3.4.4 The GIEM as the zero point
Finally, it is examined whether the halo-like excess can also be seen when the Fermi LAT GIEM model of interstellar emission [46, 47], gll_iem_v07.fits, is used as the zero point.
Though this background model is recommended for general point source analysis, it is not recommended for studying medium or large-scale structures.
The reason is that structures such as the Fermi bubbles, the GC excess, and Loop I, for which it is difficult to create reliable templates from other wavelength data, are modeled by iteratively generating non-template patches from the fit residuals.
Any components with unknown structures or shapes may be absorbed by the patch during this process, making unbiased exploration of a new diffuse component difficult.
However, the 20 GeV halo excess found in this study may still remain as an excess from the GIEM, because the spectrum of the GIEM patch is ultimately modeled by a doubly broken power-law with breaks at 150 MeV and 3 GeV. Since the 20 GeV halo excess is difficult to describe by a single power law above 3 GeV, the excess is expected to remain relative to the GIEM. In the baseline analysis of the present study, the model templates are fitted independently for each energy bin without assumptions on their spectra, while the flux distribution in the sky is fixed in a model map template. On the other hand, the GIEM patch is created to accommodate an arbitrary sky map without using templates, while imposing restrictions on the energy spectrum. Therefore, this test complements the baseline analysis.
This test is performed by a fit including the GIEM and isotropic components with non-zero initial values in MCMC (the original model normalization for the GIEM), and the NFW- model with zero initial value. However, with this method, any map component that cannot be described by a single power law spectrum will remain as an excess. To avoid confusion between the halo and other components, the other model components used in the baseline analysis (GALPROP gas and ICS, Loop I, and the structured Fermi bubble template) are also included in the fit with zero initial values in MCMC. The halo component spectrum obtained by this fit is shown in figure 14 (top). The halo component decreases compared to the baseline analysis, but is still statistically significant with the same peak at around 20 GeV (6.7, 7.4, and 5.1 at 12, 21, and 35 GeV bins, respectively). Since a part of the halo component may have been reduced by the non-template patch in the GIEM, this result can be considered as the lower limit for the halo component.
4 Implications for Dark Matter Annihilation
4.1 The possible origins of the halo-like excess
The results up to the previous section suggest that a halo-like component physically exists, with a non-power-law spectrum peaking at around 20 GeV and the morphology close to the NFW- model. The energy spectrum and morphology are consistent with what would be expected from dark matter annihilation, but other possibilities should also be considered. In principle, if there is an astrophysical source with such a spectrum and morphology, it can also explain the excess. However, no such astrophysical object or phenomenon is known, and if it exists, its origin would be an interesting topic in astrophysics.
Another possibility is that there is an unexpected, energy-dependent change in the morphology of one of the components considered in the model calculations. However, as shown in figure 13, many of the components considered have morphologies that differ greatly from a halo, and it is not easy to expect them to undergo halo-like shape changes depending on photon energy. The ICS component has the morphology closest to the spherical halo, but its shape in the GALPROP model is considerably flatter along the Galactic plane than the halo. Looking at the fit spectra for NFW- (figure 5), the halo component is less than 1/10 of the ICS at 4.3 GeV, but it is about the same or even more prominent than the ICS at 20 GeV. Explaining this by an uncertainty of the ICS model requires a model change that significantly alters the shape of the ICS map.
If adding a halo component significantly alters the ICS spectrum, rendering it unrealistic as ICS, it would suggest that a change of the ICS morphology may have been confused with a fake halo component. There is a sharp dip at the 59 GeV bin in the ICS spectrum of the NFW- fit (figure 5), but such a dip is not observed in the 21 GeV bin, where the halo component is most prominent. The 59 GeV dip is also seen, albeit shallow, in the ICS spectrum by the fit without any halo model (figure 3), and this is likely due to some anomaly in this energy bin. In other energy bins, the ICS spectrum in figure 5 is power-law-like within the range of statistical errors and is not particularly strange as an ICS spectrum. Though further investigation of the possible degeneracy between the halo and ICS components would be worthwhile in future studies, the rest of this paper discusses interpretations based on dark matter annihilation.
4.2 Fitting of the WIMP parameters
Here, the mass and velocity-averaged annihilation cross-section of the WIMPs are determined by spectral fitting, assuming that the halo excess is due to dark matter annihilation. The diffuse gamma-ray flux from annihilation per unit photon energy for an observer in a halo with a smooth density profile and no substructure (the NFW- case) is given by
where is the gamma-ray production from one annihilation event, and the integration is along the distance in the line of sight. In this study, the fitted gamma-ray flux from the halo is normalized in the Galactic pole direction (), and for the assumed MW halo model, the line-of-sight integral in this direction is
The WIMP parameters are also fitted for the case where subhalos dominate the annihilation signal, assuming that the subhalo number density is proportional to the smooth density profile (NFW-). The diffuse gamma-ray flux is expressed as
where is the annihilation rate per unit mass of dark matter, and for the assumed MW halo model,
If dark matter decays rather than annihilates, obtained by the fit to the data is related to the decay time scale . The parameter can also be linked to in the case of annihilation by the boost factor , which is widely used in the literature. This is defined as the total annihilation luminosity from all subhalos contained in a halo being boosted by a factor of , compared with that of a smooth dark matter distribution. The smallest subhalos expected from the structure formation theory are much smaller than the resolution of numerical simulations, and hence the uncertainty in is large. Estimates of hundreds or even thousands have been reported for , but more recent studies suggest 5–30 [4, 57, 58]. The combination can then be related to as
where the density integration along the Galactocentric radius for the MW halo model here is
An annihilation of dark matter is generally expected to produce a pair of a particle and its antiparticle in the Standard Model of particle physics, like . Such pairs ultimately generate numerous particles through cascading, which can be calculated using the Standard Model. Following the literature, we present three representative cases of , , and as the dominant annihilation channel, and for these are obtained from the PPPC 4 DM ID package (Release 6.0) [59, 60, 61].
Fits for the four cases (three for NFW-, and one for NFW-) are shown in figure 15, where the best-fit WIMP parameters are also shown. As photon energy decreases, statistical errors in the data become smaller, and hence spectral fitting tends to be influenced by the lowest energy bins. However, as we have already seen, the systematic error is much larger than the statistical error in these bins. Therefore, the WIMP parameters obtained here should be considered as rough estimates based on the peak position and normalization of the halo excess. For this reason, the statistical errors of the WIMP parameters are not reported. The NFW- halo flux of the baseline analysis is negative in the lowest energy bin, and the WIMP parameter fit is strongly affected by this negative value and its small statistical error. Therefore, from among the alternative analyses relative to the NFW- baseline, three are selected in which the halo flux of the lowest energy bin is not strongly negative: (1) the SNR distribution for cosmic-ray sources in GALPROP, (2) the Fermi bubble template created at 1.5 GeV, and (3) the GIEM used as the reference in the halo excess search. The case (2) results in the highest halo component flux (and therefore the largest cross section), while the case (3) results in the lowest. For the reason stated in section 3.4.4, the cross section of the case (3) should be considered as a lower limit.
In the NFW- fits, 0.4–0.8 TeV and (5–9) are favored for the and channels, and 40–80 GeV and (7–9) for the channel. The cross sections favored in the NFW- fit are ( and ) and (). These correspond to the decay timescale and s, respectively, when dark matter decays rather than annihilates. In the fits for the channel, is likely biased toward a lower value because of the small statistical errors in low-energy bins. The halo excess spectrum seems to fit better the and channels than due to their broader annihilation spectra, though the halo excess spectrum contains significant systematic uncertainties, particularly in the low-energy region.
4.3 Comparison with other dark matter searches
4.3.1 Relationship with the GC GeV excess
Dark matter annihilation is one of the leading candidates for the origin of the GC GeV excess. The 20 GeV halo excess found in this study has the peak photon energy that is significantly higher than the GC excess, and the preferred density profile ( or shallower) is flatter than that for the GC excess (). Therefore, the 20 GeV halo excess and the GC excess are likely to have different origins, and hence, if the halo excess is of dark matter origin, the GC excess must be explained by another origin. Many hypotheses have been proposed for the GC excess, and millisecond pulsars, in particular, are considered to be a hypothesis that is as promising as dark matter [8].
It should be noted that the halo excess has a flux similar to that of the GC excess at and around 2 GeV, where the GC excess peaks. If past studies have confused part of the halo excess with the extended tail of the GC excess, the true GC excess may be more centrally concentrated than . Extension of the GC excess to more than 10∘ has been argued to support the dark matter origin [62, 63, 64], but if the true GC excess is more centrally concentrated, it may be more consistent with stellar population origins, such as millisecond pulsars.
4.3.2 Comparison with previous constraints and theoretical expectations
The annihilation cross section estimated from the 20 GeV halo excess ( with GeV for the channel) is marginally consistent with the upper limits obtained by earlier studies on the MW halo region [26, 27, 28, 29, 31]. The LAT team’s analysis using two years of the LAT data [26] reported without background modeling, for their standard choice of the dark matter density at the solar neighborhood . It should also be noted that the ROI of the LAT team’s analysis (, ) is much narrower in the Galactic latitude than that of this study.
The cross section required for the halo excess is larger than the upper limits by dwarf spheroidal galaxies, which are currently considered to be the strongest [10, 11, 12, 13]. For example, it is several times larger than the 95% containment region of the constraints obtained in ref. [13]. However, this tension does not immediately rule out the dark matter interpretation of the 20 GeV halo excess. Compared to dwarf galaxies, whose dark matter mass can be more reliably constrained by observations, the density distribution of the MW halo is highly uncertain ([4], see also section 4.3.3 below). The dark matter content in dwarf galaxies is also still subject to uncertainties regarding methodology and observed data [65, 66]. In fact, the dark matter interpretation of the GC GeV excess is also in the same degree of tension with dwarf galaxy constraints, but for the same reason, the dark matter interpretation is not considered to be excluded [8, 13].
The annihilation cross section determines the era of WIMP decoupling in the early universe, and is therefore related to the relic amount of dark matter in the present universe. The -channel cross section estimated from the 20 GeV halo excess is more than an order of magnitude larger than the canonical thermal relic cross section to explain the present dark matter density, (2–3) [6, 4]. However, considering the uncertainty and degrees of freedom in particle physics theory, there are various possibilities that the actual annihilation cross section differs from this by orders of magnitude. Considering also the uncertainty in the MW halo density profile, the dark matter interpretation of the 20 GeV halo excess is not inconsistent with theoretical expectations.
4.3.3 Uncertainties in the MW dark matter density distribution
The uncertainties in the density profile of the MW halo can be described in more detail as follows. Since the mass of baryons exceeds that of dark matter in the inner region of the MW ( 10 kpc) [67, 68], we can only extrapolate from the dark matter density at the solar system () assuming a halo density profile, but is uncertain by a factor of a few. There are several other analytic functions proposed for the halo density profile, in addition to NFW, and the Einasto profile [69], which is known to fit numerical simulations as well as NFW, is particularly interesting for the halo excess. When is fixed, this profile predicts a higher density than NFW at around 10∘ ( kpc) from the GC, and a shallower profile than NFW near the GC (see figure 2.3 of [4]), which are in agreement with the trends suggested by the 20 GeV halo excess.
In addition to the uncertainty in the fitting formulae for the mean halo density profile in numerical simulations, individual halo profiles also vary depending on their formation and merger history. Furthermore, in the central region of a halo, the dark matter density profile may be affected by complicated baryon dynamics, and this effect is not yet fully understood [70, 71]. In the case of a smooth density distribution (), a threefold uncertainty in the estimate of dark matter density in the 10∘–20∘ region from the GC changes the halo flux by nearly an order of magnitude. Finally, the effect of subhalos or any substructures in the MW halo should not be overlooked. Since substructures are more easily destroyed in the central region [52], the signal may be boosted with an effective density profile shallower than NFW, in a direction consistent with the trends of the halo excess.
4.3.4 The extragalactic gamma-ray background radiation
If the MW halo emits annihilation gamma-rays, the extragalactic gamma-ray background radiation (EGB) should include a contribution from annihilations in all dark halos in the universe. Theoretical predictions suggest that the EGB flux by annihilation is comparable to that from the MW halo, with the peak photon energy redshifted to about half [22]. The observed EGB in the 1–1000 GeV range is thought to contain various astronomical components, such as active galactic nuclei and star-forming galaxies, making it difficult to identify the dark matter component unambiguously. A recent study suggests that a significant contribution to the EGB from annihilation with the WIMP parameters of the 20 GeV halo excess is not ruled out [25].
The EGB from star-forming galaxies is well understood in terms of physical processes, and relatively reliable theoretical models have been developed [23, 72]. These models contribute significantly to the Fermi LAT data below 10 GeV, but their predicted flux decreases rapidly above 10 GeV due to intergalactic absorption. On the other hand, the EGB data extend approximately as a power law beyond 10 GeV, and a new component may be appearing in this region. Annihilation gamma rays from the MW halo and extragalactic halos, whose spectral peak is around 10–20 GeV, may be a candidate for such a component. In fact, this was the original motivation behind this study to search for signs of dark matter annihilation.
5 Summary and Discussion
In this study, 15 years of the Fermi LAT data have been used to search for annihilation gamma rays from dark matter in the halo region of the MW (, ), in the photon energy range of 1.5–810 GeV. Following the standard analyses by the LAT team, the gamma-ray intensity map in the ROI is modeled considering the following components: point sources, the interstellar gas and ICS components of GALPROP, isotropic background radiation, Loop I, and the Fermi bubbles. Then it is investigated whether a halo-like component is necessary in addition to these model components, in the ROI excluding the Galactic plane region (). To focus on the smooth structure of the halo scale, photon counts for the likelihood are calculated in each analysis cell of size .
The model fitting analysis favors the presence of a halo component with a spectral peak around 20 GeV at high statistical significance, in contrast to the GC GeV excess that peaks at 2–3 GeV. The flux of the halo component is consistent with zero at energies below 2 GeV and above 200 GeV. The radial profile fits well with the NFW- model (the NFW profile and gamma-ray emissivity per unit volume for the smooth distribution, ), rather than (GC GeV excess) or (subhalos or decaying dark matter). To avoid exceeding the data when extrapolating the halo component to the GC, a profile that is slightly shallower than NFW- is preferred, especially in the central region. Looking at the maps of the halo model plus fit residual, the shapes are nearly spherically symmetric in all energy bins where the halo component is detected, indicating that the halo model is a good description of the observed map.
Various types of systematic uncertainties are examined, including the GALPROP model parameters, the energy bin used to create the Fermi bubble template, and the morphology of the bubble template. Due to the strong diffuse emission from cosmic-ray interactions, the systematic uncertainty of the halo component flux becomes larger at lower photon energies. However, the trend that the halo component rises from 1 GeV, peaks at around 20 GeV, and becomes almost undetectable above 200 GeV remains unchanged. In particular, it is noteworthy that the halo-like excess is detected with a similar spectrum, even when the LAT standard background model (GIEM) is used as the zero point. Since the GIEM includes a non-template patch created from the fit residuals, this suggests that the halo excess is not an artifact dependent on the model map templates used in the baseline analysis.
While the possibility of astrophysical explanation cannot be ruled out, the characteristics of this “Fermi halo” are consistent with what would be expected for dark matter annihilation signals. The spectrum of the halo excess can be fitted with the most representative annihilation channel into , and in that case, the mass and annihilation cross section of the WIMPs are estimated to be 0.5–0.8 TeV and (5–8), respectively. This cross section is in tension with the upper limits derived from dwarf spheroidal galaxies, and is more than an order of magnitude larger than the canonical thermal relic cross section. However, considering various uncertainties, particularly those in the density profile of the MW halo, the dark matter interpretation of the Fermi halo is compatible with these.
However, the signal from the MW halo alone does not constitute the definitive proof of dark matter annihilation. Detection of annihilation signals from other objects or regions with consistent WIMP parameters will be crucial for the final confirmation. Gamma-ray observations of dwarf galaxies in the MW halo are particularly interesting from this perspective. Given the tension with the constraints from dwarf galaxies, a modest improvement in the sensitivity of future observations should eventually lead to the detection of annihilation gamma rays. It is noteworthy that interesting excesses have already been observed in several dwarf galaxies, even if they have not been definitively detected [13]. The largest excess has been reported in Reticulum II by several studies (typically –) [73, 74, 75, 76, 13], and the WIMP mass estimated by the latest study [13] is interestingly similar to that for the MW halo excess found by the present study.
Another important avenue is the search for line gamma-ray emission. Line emissions at the photon energy equal to the WIMP mass (0.3–0.8 TeV for the channel) are expected, although the intensity will depend on the branching ratio of annihilation into two photons (or perhaps ). In this energy range, ground-based atmospheric Cherenkov telescopes are more sensitive than the Fermi LAT, and further significant improvements in sensitivity are expected shortly by the Cherenkov Telescope Array Observatory (CTAO) [77].
Dark matter search using neutrinos is also interesting. Though the current neutrino constraints [4, 78, 79, 80] are 1–2 orders of magnitude (depending on the annihilation channels) weaker than the mass and cross section suggested in this work, future improvements in the sensitivity of neutrino detectors may enable independent verification. Since the gamma-ray to neutrino flux ratio can be calculated using the Standard Model of particle physics, the neutrino flux from the Fermi halo can be predicted without uncertainty about the dark matter density profile.
Acknowledgments
The author thanks Tsunefumi Mizuno for useful discussion and advice. The author was supported by the JSPS/MEXT KAKENHI Grant Number 18K03692.

### v2

@@
 Gamma-ray maps within the region of interest (, ) are modeled using point sources, the GALPROP models of cosmic-ray interactions, isotropic background, and templates of Loop I and the Fermi bubbles, and then the presence of a halo-like component is further examined.
-A statistically significant halo-like excess is found with a sharp peak around 20 GeV, while its flux is consistent with zero below 2 GeV and above 200 GeV.
+A statistically significant halo-like excess is found with a spectral peak around 20 GeV, while its flux is consistent with zero below 2 GeV and above 200 GeV.
 Examination of the fit residual maps indicates that a spherically symmetric halo component fits the map data well.
@@
 Various systematic uncertainties are investigated, but the 20 GeV peak remains significant.
-In particular, the halo excess with a similar spectrum is detected even relative to the LAT standard background model, which does not depend on GALPROP or other model templates.
-The halo excess can be fitted by the annihilation spectrum with a mass 0.5–0.8 TeV and annihilation cross section (5–8) for the channel.
+In particular, the halo excess with a similar spectrum is detected even relative to the LAT standard background model, which contains non-template patches adjusted to match the observed map.
+The halo excess spectrum can be fitted by annihilation with a particle mass 0.5–0.8 TeV and cross section (5–8) for the channel.
 This cross section is larger than the upper limits from dwarf galaxies and the canonical thermal relic value, but considering various uncertainties, especially the density profile of the MW halo, the dark matter interpretation of the 20 GeV “Fermi halo” remains feasible.
@@
 Even though nearly a century has passed since Fritz Zwicky inferred the existence of dark matter from the motion of galaxies in the Coma cluster [1], dark matter remains one of the greatest mysteries in astrophysics, cosmology, and fundamental physics [2, 3, 4].
-Weakly interacting massive particles (WIMPs) in the mass range of 1 GeV to 100 TeV have long been considered one of the leading candidates, because they interact only through the weak interaction, and because the thermal relic amount with the annihilation cross section expected from particle physics is close to the observed dark matter density [3, 4, 5, 6].
+Weakly interacting massive particles (WIMPs) in a mass range from 1 GeV to 100 TeV near the electroweak scale have long been considered one of the leading candidates, because such particles are naturally predicted in some particle physics theories beyond the Standard Model, and because the thermal relic amount predicted from the annihilation cross section expected for WIMPs is close to the observed dark matter density [3, 4, 5, 6].
 The existence of WIMPs can be probed with direct searches by underground experiments, indirect searches using annihilation products (e.g., gamma-rays and neutrinos), and by particle accelerator experiments.
-Regarding indirect searches by gamma-rays, the sensitivities of the Fermi Large Area Telescope (Fermi LAT) and ground-based air Cherenkov gamma-ray telescopes have reached a level where the intensity expected from the canonical annihilation cross section can be detected.
-When WIMPs directly annihilate into two photons, the line gamma-ray emission is easy to distinguish, but the branching ratio into line emission is highly uncertain.
+Regarding indirect searches by gamma rays, the sensitivities of the Fermi Large Area Telescope (Fermi LAT) and ground-based air Cherenkov gamma-ray telescopes have reached a level where the intensity expected from the canonical annihilation cross section can be detected.
+When WIMPs directly annihilate into two photons, line (monoenergetic) gamma-ray emission is easy to distinguish, but the branching ratio into line emission is highly uncertain.
 In the more mainstream annihilation mode, the Standard Model particles such as or are first generated, followed by a cascade that produces gamma rays with a continuum energy spectrum.
-The spectrum still shows a broad peak around a specific energy related to the WIMP mass, which contrasts with power-law spectra of many astrophysical objects.
-Numerous explorations have been conducted toward various astronomical objects or regions in search of such signals, including the Milky Way (MW) Galactic center (GC) [7, 8, 9], dwarf galaxies [10, 11, 12, 13] and dark subhalos [14, 15] in the MW halo, the Andromeda galaxy [16, 17], galaxy groups and clusters [18, 19, 20, 21], and the extragalactic background radiation [22, 23, 24, 25].
+The spectrum still shows a broad peak around a specific gamma-ray energy related to the WIMP mass, which contrasts with power-law spectra of many astrophysical objects.
+Numerous explorations have been conducted toward various astronomical objects or regions in search of such signals, including the Milky Way (MW) Galactic center (GC) [7, 8, 9], the Large and Small Magellanic Clouds (LMC and SMC) [10, 11], dwarf galaxies [12, 13, 14, 15] and dark subhalos [16, 17] in the MW halo, the Andromeda galaxy [18, 19], galaxy groups and clusters [20, 21, 22, 23], and the extragalactic background radiation [24, 25, 26, 27].
 This study will search for annihilation gamma rays from the MW halo, avoiding the GC and disk regions.
@@
 Then the true dark matter signal may appear in other photon energy ranges.
-The dark matter interpretation of the GC excess suggests that the radial density profile near the GC is with 1.1-1.3, which is more centrally concentrated than the standard Navarro-Frenk-White (NFW) [38] profile ( in the central region).
-If the true density profile is the NFW or less centrally concentrated, dark matter signals may be more easily detected in the halo region than at the GC.
-In 2012, the Fermi LAT team analyzed the MW halo region in search of dark matter using two years of the LAT data, which is much smaller than the amount accumulated to date.
+To explain the GC excess with dark matter annihilation, the radial density profile near the GC must be with 1.1–1.3, which is more centrally concentrated than the standard Navarro-Frenk-White (NFW) [40] profile ( in the central region).
+If the true density profile is NFW or less centrally concentrated, dark matter signals may be more easily detected in the halo region than at the GC.
+In 2012, the Fermi LAT team analyzed the MW halo region in search of dark matter using two years of the LAT data [28], which is much smaller than the amount accumulated to date.
 From that time until now, understanding and analysis methods for diffuse radiation in the halo region, such as the Fermi bubbles and Loop I, have also matured.
@@
 2.2 Likelihood analyses
-In many studies fitting a model map to the LAT data, the photon counts and the model expectations are compared in each pixel.
-However, the halo-like component is expected to be much smoother than the pixel scale of 0.125∘.
+The maximum likelihood fitting is performed independently for each photon energy bin.
+In many studies that fit a model map to the LAT data, the photon counts and the model expectations are compared in each pixel.
+Since the halo-like component is expected to be much smoother than the pixel scale of 0.125∘, calculating the likelihood at a coarser resolution allows computation time to be reduced without compromising dark matter search.
 Therefore, we divide the ROI into cells with a width of and calculate the likelihood based on photon counts and expectations in these cells.
-This not only speeds up calculations, but also prevents the likelihood from being unnecessarily biased by pixel-scale fine discrepancies between observed counts and model predictions.
-The maximum likelihood fitting is performed independently for each photon energy bin.
+Some components of the Galactic diffuse emission, such as those from interstellar gas, possess pixel-scale fine structures, which may help separate them from others by pixel-scale likelihood calculation.
+On the other hand, smoothing at coarse resolutions may reduce the impact of pixel-scale mismatches between a model map and data on the likelihood estimate.
 Let be the observed photon count in the -th cell and be the number of photons expected from the model, and then the likelihood function based on Poisson statistics is
@@
 4FGL-DR4 also provides a catalog of extended sources, in addition to the point sources.
-Since the extended sources are not many, circular regions with a radius twice that of the semimajor axis of the catalog are masked and excluded from the analysis.
+Since the extended sources are not many, circular regions with a radius twice the semi-major axis of the catalog are masked and excluded from the analysis.
 Gamma rays produced by cosmic rays propagating in the MW are the most significant component of the Galactic diffuse gamma-ray emission.
@@
 Among the various GALPROP parameter sets used in the LAT team’s study of the Galactic diffuse emission [47], SLZ6R30T150C2 is adopted as the baseline model in the present work.
-This parameter set is used for the inverse Compton scattering (ICS) component in the Galactic interstellar emission model (GIEM, gll_iem_v07.fits)555https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html [46, 47] recommended by the LAT team for normal point source analysis.
+This parameter set is used for the inverse Compton scattering (ICS) component in the Galactic interstellar emission model (GIEM, gll_iem_v07.fits)555https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html [48, 49], which is recommended by the LAT team for normal point source analysis.
 The GALPROP model radiation is divided into two components, gas and ICS, which are fitted independently.
 The gas component is the sum of pion-decay gamma rays produced by cosmic-ray hadrons, and bremsstrahlung from cosmic-ray electrons and positrons.
-The GALPROP model emission can be further divided into more components, and the fitting results with such a larger number of sub-components or different model parameters will be discussed in section 3.4.3.
+The flux ratio between pion decay and bremsstrahlung is fixed to the original model value, making the gas component a single template.
+The GALPROP model emission can be further divided into different components, and the fitting results with a larger number of sub-components or different model parameters will be discussed in section 3.4.4.
 Loop I is a giant structure spanning more than 100∘ in the radio sky [50], and although it is unclear whether it is nearby supernova remnants or an outflow from the GC [51, 52], it has also been detected by LAT [53].
@@
 The parameters of the geometric model are taken from the LAT team papers [39, 7], which differ slightly from the values in Wollben’s original paper [51].
-The normalization factors for each of the two shells are treated as independent fit parameters ().
+The normalization factors for each of the two shells are treated as two independent fit parameters ().
 To calculate the halo component model map, we first assume that the density profile of the MW dark matter halo is NFW, and use the parameters from the Via Lactea II simulation [54, 55].
 Therefore, the density profile within the virial radius = 402 kpc is , where , 21 kpc, and .
-The distance from the GC to the solar system is set at 8 kpc.
+The distance from the GC to the solar system is set at 8 kpc, and hence .
 Annihilation gamma-ray diffuse flux is proportional to the line-of-sight integration of the gamma-ray emissivity per unit volume.
@@
 Therefore, the map template of (NFW-) will also be tested.
-Subhalos may be destroyed by strong tidal forces near the center of the parent halo [52], in which case emissivity in the central region would be suppressed, but this is not tested in this study.
-Finally, based on previous studies indicating that the density profile of () is preferred for the GC GeV excess, the model template of (NFW-), which becomes the same emissivity profile as the GC excess in the central region using the original NFW, is included in the fit for this component.
+Subhalos may be destroyed by strong tidal forces near the center of the parent halo [54], in which case emissivity in the central region would be suppressed.
+While a model for this case is not tested in this work, this possibility should be kept in mind.
+Previous studies on the GC GeV excess found that the density profile of () is preferred, and hence the template of with the original NFW (NFW-) is adopted as the model of the GC excess, whose emissivity becomes in the central region ().
 The map morphologies of the point sources and GALPROP have photon energy dependencies determined by the models, while other model maps do not depend on photon energy.
 The initial values of in the MCMC chain are the same as the original normalization for the models of point sources and GALPROP.
-The initial of the isotropic component is set to , which is typical of the reported fluxes of the extragalactic gamma-ray background radiation [54].
+The initial of the isotropic component is set to , which is typical of the reported fluxes of the unresolved extragalactic gamma-ray background radiation [56].
 For other components, the initial values of the MCMC chain are set to zero.
@@
 Therefore, the map template for the bubbles used in this study is made here, mostly following the methods adopted in the LAT team’s papers [39, 7].
-As the first step, a flat (i.e., isotropic) template within the rough area of the Fermi bubbles is assumed, and it is fit to the data together with the other model components.
+Two photon energy bins, 1.5 and 4.3 GeV, are selected for this procedure, which are in the typical photon energy range where the Fermi bubbles are prominent.
+As the first step, a flat (i.e., isotropic) template with sharp edges within a rough area of the Fermi bubbles is assumed, and it is fit to the data together with the other model components.
 Here, the GC GeV excess is included in the fit as a known component, but other halo components are not considered.
 At this stage, the fit is performed in the whole ROI, including the Galactic disk region (), to ensure an accurate fit to the GC excess.
-Looking at the best-fit image of the bubbles (i.e., sum of the flat template and fit residuals, smoothed by the Gaussian with ), the energy-independent boundary of the flat template is iteratively improved.
-The image of the Fermi bubbles obtained in this way is shown in figure 1, together with the final boundary of the flat template, in the two energy bins of 1.5 and 4.3 GeV, which are in the typical photon energy range where the Fermi bubbles are prominent.
+The best-fit image of the bubble is created from the sum of the flat template and fit residuals; the photon count map of this is converted into physical flux at each pixel using the exposure data, and then smoothed with a Gaussian filter with .
+The energy-independent boundary of the flat template is iteratively improved by looking at these best-fit images.
+The image of the Fermi bubbles obtained in this way in the two energy bins is shown in figure 1, together with the final boundary of the flat template.
 Since the template of the Fermi bubbles will only be used at in subsequent halo component searches, the Galactic disk region () is not shown.
@@
 The fit in this study is performed on a cell-by-cell basis with the cell size of , and PSF errors should have little effect on the results.
-The best-fit spectra (calculated by the best-fit in each energy bin) for all the model components by this fit are shown in figure 2.
-Here, it should be noted that the spectral flux shown in this figure excludes the disk region for comparison with the later analyses, though the model fitting is performed including the disk region.
-The spectra of the model templates are generally consistent with the results of similar analyses by the LAT team [7], indicating that there are no serious problems or errors in this analysis.
-In particular, the peak photon energy and flux of the GC GeV excess ( at 1∘ from the GC at 2.5 GeV) are also consistent with those reported in the literature [8].
+The best-fit spectra of all the model components by the fit with the final flat bubble template are presented in figure 2.
+Here, the mean background flux within the ROI, including the Galactic disk region, is shown.
+For comparison with later analysis, figure 3 shows the flux excluding the disk, though the model fitting is the same as figure 2, i.e., including the disk.
+The spectra of the model templates are generally consistent with those of similar analyses by the LAT team [7], indicating that there are no serious problems or errors in this analysis.
+In particular, the peak photon energy and the flux of the GC GeV excess ( at 1∘ from the GC at 2.5 GeV) are also consistent with those reported in the literature [8].
 The morphology of the Fermi bubbles is also consistent with what has been reported in the literature.
@@
 This feature corresponds to the regions of particularly strong flux in the gas component map of GALPROP (see figure 14 in section 3.3 later), and is likely due to discrepancies between the GALPROP model and the data.
-A similar feature is seen in the LAT team’s analysis using GALPROP (Fig.
-22 of ref.
+A similar feature is seen in the LAT team’s analysis using GALPROP (figure 22 of ref.
 [39]), and the LAT team used the method of spectral component analysis to produce a bubble template without this feature.
 Since the purpose of the present study is not to create an accurate template of the Fermi bubbles, the following approach will be taken to explore an additional halo component.
-Among the two energy bins, the 4.3 GeV map is chosen as the source to create the template, because the Fermi bubbles are more clearly seen.
+Among the two energy bins, the 4.3 GeV map is chosen as the source to create the structured bubble template, because the Fermi bubbles are more clearly seen.
 In the map shown in figure 1, the positive flux regions are dominated by the Fermi bubbles.
 Therefore, the map of positive regions, regardless of whether it is inside or outside the boundaries of the flat template, will be used as the energy-independent template for the bubbles.
-On the other hand, in order to eliminate the effects of discrepancies between GALPROP and the data, the map of the negative flux regions is introduced into the fit as another template.
+On the other hand, to minimize the impact of the discrepancy between GALPROP and the data, the map of the negative flux regions is introduced into the fit as another template.
 Since these two probably have different energy spectra, they are included as two independent templates.
+Hereafter, they will be referred to as the positive and negative residual templates, though these strictly include not only the fit residual but also the flat bubble template.
 With this treatment, the model and data maps should agree well without additional model components at around 4.3 GeV.
@@
 This is the signal we will be searching for.
-3.2 Fitting results with a halo-like component
+3.2 Fitting with a halo-like component
 From here, the fits will be performed excluding the Galactic disk region () to search for a halo-like component.
 The flat template of the Fermi bubbles is replaced with the positive and negative residual templates obtained in the previous section.
-The fit parameter is limited to for the positive residual template (the Fermi bubbles), but not for the negative template, because the inverted negative template (i.e., positive flux) may be favored for a better fit, if the true morphology of the Fermi bubbles is relatively flat and the discrepancy between GALPROP and the data decreases in higher energy bins,
+The fit parameter is limited to for the positive residual template (the Fermi bubbles), but not for the negative template, because the inverted negative template (i.e., positive flux) may be favored for a better fit, if the true morphology of the Fermi bubbles is relatively flat and the discrepancy between GALPROP and the data decreases in higher energy bins.
 First, the fit is performed without the GC excess or any other halo component, and the best-fit spectra are shown in figure 4.
@@
 This means that extrapolating this fit to the GC would significantly exceed the observed data.
-These results indicate that a halo-like component exists around 20 GeV, with a density profile shallower than that assumed for the GC GeV excess, and is prominent in the region.
-Therefore, to test a flatter density profile, we perform fits using the NFW- and NFW- halo models instead of the GC excess, and the results are shown in figures 5 and 6, respectively.
+These results imply that a halo-like component exists around 20 GeV and is prominent in the region, with a density profile shallower than that assumed for the GC GeV excess.
+Therefore, to test a flatter density profile, we perform fits using the NFW- and NFW- halo models instead of the GC excess, and the results are shown in figures 6 and 7, respectively.
 Only one of the three halo models is used in a fit to avoid degeneracy between the similar halo maps.
@@
 Another possibility is that part of the halo component at 4.3 GeV has been absorbed in the residual map templates.
-If the halo component weakens toward lower energy below 4.3 GeV, the fit of the halo component could be negative to cancel out the halo component absorbed in the residual map templates.
+If the halo component weakens toward lower energy below 4.3 GeV, the fit of the halo component could be negative to cancel out the part absorbed in the residual map templates.
 Figure 9 illustrates the likelihood (relative to the no-halo fit) of the fits with three different halo models, to compare the goodness of fit.
 In the 2.5 GeV bin, the halo components are fitted to be nearly zero (figure 8), and a deep dip is seen in the figure.
-Around the peak photon energy of the halo excess (10–20 GeV), the flatter NFW- model fits better than NFW- at 95% confidence level (CL).
-On the other hand, the shallowest NFW- model fits even worse than the NFW- model (at a significance much higher than 95% CL), suggesting that a profile close to NFW- best fits the data in the region of .
+Around the peak photon energy of the halo excess (10–20 GeV), the NFW- model fits better than the steepest NFW- at 95% confidence level (CL).
+On the other hand, the shallowest NFW- model fits even worse than the NFW- (at a significance much higher than 95% CL), suggesting that a profile close to NFW- best fits the data in the region of .
 3.3 Morphology and radial profile of the halo-like excess
-Figure 9 shows the radial angular profiles of three halo model fits, at the energy bin of 21 GeV where the halo component is the most prominent.
+Figure 10 shows the radial angular profiles of three halo model fits at the energy bin of 21 GeV, where the halo component is the most prominent.
 The diffuse flux of the best-fit halo model is shown as a curve, and the data points are the sum of the best-fit model and the fit residuals, in a ring region centered on the GC within the ROI (, ).
-This plot confirms what we found in the previous section, namely that the NFW- model best fits the data in terms of likelihood.
+It is not straightforward to determine from this figure which of the three halo models fits best, but as seen in the previous section 3.2 and figure 9, the likelihood values indicate that the NFW- model is preferred with a statistical significance of 2 or greater.
 The density profile of the NFW- model is steeper than the data, while the NFW- model is shallower.
-For comparison, the best fit flux of the GC GeV excess (the NFW- model in the fit including the Galactic plane) is also shown in the figure as the curve of “NFW- ()”.
+The data points differ significantly among the three models, especially in the tail region, suggesting the presence of degeneracy with other model components.
+In particular, the tail of the halo models is considered prone to degeneracy with the isotropic component.
+Indeed, in the NFW- fit, the isotropic component decreases significantly near 20 GeV compared to the NFW- fit, becoming consistent with zero (figures 6 and 7).
+In the NFW- fit, the flux of the isotropic component near 20 GeV (figure 6) agrees well with the unresolved isotropic background flux reported by the Fermi-LAT team [56], (shown in figure 10).
+This implies that the NFW- fit leaves no room for the isotropic background that should inherently exist, providing further evidence that the NFW- model fits the data better, in addition to the likelihood values.
+For comparison, the best fit flux of the GC GeV excess (the NFW- model in the fit including the Galactic plane, figure 2) is also shown in the figure as the curve of “NFW- ()”.
 The halo component must have a shallow profile that does not exceed this curve when extrapolated to the GC.
@@
 When the fit is made without a halo model component, the fit tries to minimize the halo component, and hence, the fit would be biased if the halo component physically exists.
-However, a halo-like residual can be seen even in such a fit (top left).
+However, a halo-like residual can be seen even in such a fit (top left), giving a lower bound for the halo-like signal.
 In the fit including NFW- as the halo model, the map of the model plus residual (i.e., the observed map with all model maps except for the halo subtracted) shows a clearer halo-like structure (top right), which is spherically symmetric and consistent with the model map (bottom left).
 The fit residual with the halo model (bottom right) is significantly reduced compared to the fit without the halo (top left), indicating that the halo fit is successful overall in the ROI.
-To examine the fits in energy bins other than the peak (21 GeV), figures 11 and 12 show maps of the NFW- model plus residual at bins lower and higher than 21 GeV, respectively.
+To examine the fits in energy bins other than the peak (21 GeV), figures 12 and 13 show maps of the NFW- model plus residual at bins lower and higher than 21 GeV, respectively.
 These maps show the trend seen in figure 8, with the negative fit at the lowest 1.5 GeV, nearly zero at 2.5 GeV, the halo component becoming stronger toward 21 GeV, and then rapidly weaker beyond 21 GeV.
-The morphology of the halo model plus residual is generally spherically symmetric and does not appear to contradict the profile of the assumed model halo.
+The morphologies of the halo model plus residual are generally spherically symmetric and do not appear to contradict the profile of the assumed model halo.
 Finally, the images of the best-fit model map templates considered in the fits, other than the halo, are shown in figure 14 for comparison.
@@
 3.4 Examination of systematic uncertainties
-Here, we examine systematic uncertainties concerning the evidence for the halo-like excess on the following points: (1) uncertainties in the Fermi bubble template, (2) treatment of point sources, (3) model parameters and number of free components in GALPROP, and (4) halo-like excess search using the Fermi LAT standard background model (GIEM) as the reference.
+Here, systematic uncertainties concerning the evidence for the halo-like excess are examined on the following points: (1) fit analysis divided into four quadrants, (2) uncertainties in the Fermi bubble template, (3) treatment of point sources, (4) model parameters and number of free components in GALPROP, and (5) halo-like excess search using the Fermi LAT standard background model (GIEM) as the reference.
+3.4.1 Analysis divided into four quadrants of the Galactic coordinates
+To test the isotropy of the halo excess around the GC, the NFW- fits are performed by dividing into four quadrants about the Galactic coordinates, and the halo excess spectra are shown in figure 15 (top panel).
+Furthermore, the fitting results are also illustrated when dividing the four quadrants shifted by 45∘ into two quadrants in the galactic plane direction () and two quadrants in the polar direction ().
+Although statistical fluctuations increase significantly above 30 GeV due to the decrease in photon number, a consistent trend is observed in all quadrants below 20 GeV.
 3.4.2 The Fermi bubble template
 The Fermi bubble template was created from the positive residual map at 4.3 GeV; however, a portion of the halo component in that energy bin may be absorbed by the template, resulting in an underestimate of the halo component in the fit.
-To check the dependence on the choice of energy bin in this process, an alternative fit result is shown in Figure 14 (top), in which the positive (the Fermi bubbles) and negative residual templates are replaced with those of the 1.5 GeV map shown in figure 1.
+To check the dependence on the choice of energy bin in this process, an alternative fit result is shown in figure 15 (middle), in which the positive (the Fermi bubbles) and negative residual templates are replaced with those of the 1.5 GeV map shown in figure 1.
 As expected, the energy bin where the halo component becomes zero has shifted to the lower side, and the halo flux at the 20 GeV peak has increased, but there is no change in the basic shape of the spectrum with a peak near 20 GeV.
 As shown in figure 1, due to the strong negative residual around 10–15∘, the positive residual template may be biased from the true structure of the Fermi bubbles.
-To test the dependence on the Fermi bubble template structure, the fit result obtained by replacing the positive residual template with the simple flat template is shown in Figure 14 (top).
-The halo component has decreased compared to the baseline analysis, and this may be because the baseline analysis attempted to compensate for the concavity caused by the negative residual with the halo component.
+To test the dependence on the Fermi bubble template structure, the fit result obtained by replacing the positive residual template with the simple flat template is shown in figure 15 (middle).
+The halo component has decreased compared to the baseline analysis, likely because the bubble flux in the negative residual region becomes larger compared to the structured bubble template, and the halo component is fitted smaller to compensate for this.
 However, the overall spectral shape, including the rising rate of the halo component from 1 to 20 GeV, remains unchanged.
 The excess at the 21 GeV bin is still statistically significant (9.1).
+As a further verification of systematic uncertainties related to the negative residual region, the fit is performed after this region is masked.
+Pixels with notably negative flux () in the right panel of figure 1 are excluded from the ROI, and the negative residual map is removed from the set of the fitting templates.
+The fit result is shown in figure 15 (middle); the halo excess spectrum does not differ significantly from the baseline analysis.
 3.4.3 Point sources
 In the baseline analysis, point sources are subtracted using the spectral models in the 4FGL-DR4 catalog and the LAT PSF.
-Instead, the fit result after masking circular areas around the point sources with a radius equal to the 95% containment angle is shown in figure 14 (top).
+Instead, the fit result after masking circular areas around the point sources with a radius equal to the 95% containment angle is shown in figure 15 (middle).
 The result is almost unchanged from the baseline analysis.
@@
 To check the dependence on the number of free GALPROP components, the NFW- fit is performed with the same treatment about GALPROP as ref.
-[7], and the result is shown in Figure 14 (top).
+[7], and the result is shown in figure 15 (middle).
 The halo component decreases in bins below 12 GeV from the baseline analysis, but remains almost unchanged at the 21 GeV bin, and the overall trend of the spectrum peaking at 20 GeV remains unchanged.
@@
 Though this background model is recommended for general point source analysis, it is not recommended for studying medium or large-scale structures.
-The reason is that structures such as the Fermi bubbles, the GC excess, and Loop I, for which it is difficult to create reliable templates from other wavelength data, are modeled by iteratively generating non-template patches from the fit residuals.
+The reason is that diffuse structures such as the Fermi bubbles, the GC excess, and Loop I, for which it is difficult to create reliable templates from other wavelength data, are modeled by iteratively generating non-template patches from the fit residuals.
 Any components with unknown structures or shapes may be absorbed by the patch during this process, making unbiased exploration of a new diffuse component difficult.
@@
 Therefore, this test complements the baseline analysis.
-This test is performed by a fit including the GIEM and isotropic components with non-zero initial values in MCMC (the original model normalization for the GIEM), and the NFW- model with zero initial value.
-However, with this method, any map component that cannot be described by a single power law spectrum will remain as an excess.
-To avoid confusion between the halo and other components, the other model components used in the baseline analysis (GALPROP gas and ICS, Loop I, and the structured Fermi bubble template) are also included in the fit with zero initial values in MCMC.
-The halo component spectrum obtained by this fit is shown in figure 14 (top).
+This test is performed by a fit that includes the GIEM, point sources, and the isotropic background with non-zero initial values in MCMC (the original model normalization for the GIEM), as well as the NFW- model with a zero initial value.
+However, with this treatment, any map component that a simple power law spectrum cannot describe will remain as an excess.
+To avoid confusion between the halo and other components, the model components used in the baseline analysis (GALPROP gas and ICS, Loop I, and the structured Fermi bubble template) are also included in the fit with zero initial values in MCMC.
+The halo component spectrum obtained by this fit is shown in figure 15 (middle).
 The halo component decreases compared to the baseline analysis, but is still statistically significant with the same peak at around 20 GeV (6.7, 7.4, and 5.1 at 12, 21, and 35 GeV bins, respectively).
-Since a part of the halo component may have been reduced by the non-template patch in the GIEM, this result can be considered as the lower limit for the halo component.
+Since a part of the halo component may have been reduced by the non-template patch in the GIEM, this result can be considered as a lower limit for the halo component.
 4 Implications for Dark Matter Annihilation
@@
 For this reason, the statistical errors of the WIMP parameters are not reported.
-The NFW- halo flux of the baseline analysis is negative in the lowest energy bin, and the WIMP parameter fit is strongly affected by this negative value and its small statistical error.
+The NFW- halo flux of the baseline analysis is negative in the lowest energy bin, and the WIMP parameter fit is strongly affected by this unphysical negative value and its small statistical error.
 Therefore, from among the alternative analyses relative to the NFW- baseline, three are selected in which the halo flux of the lowest energy bin is not strongly negative: (1) the SNR distribution for cosmic-ray sources in GALPROP, (2) the Fermi bubble template created at 1.5 GeV, and (3) the GIEM used as the reference in the halo excess search.
@@
 In the fits for the channel, is likely biased toward a lower value because of the small statistical errors in low-energy bins.
-The halo excess spectrum seems to fit better the and channels than due to their broader annihilation spectra, though the halo excess spectrum contains significant systematic uncertainties, particularly in the low-energy region.
+The halo excess spectrum seems to fit the and channels better than due to their broader annihilation spectra, though the halo excess spectrum contains significant systematic uncertainties, particularly in the low-energy region.
 4.3 Comparison with other dark matter searches
@@
 Many hypotheses have been proposed for the GC excess, and millisecond pulsars, in particular, are considered to be a hypothesis that is as promising as dark matter [8].
-It should be noted that the halo excess has a flux similar to that of the GC excess at and around 2 GeV, where the GC excess peaks.
+It should be noted that the halo excess has a flux similar to that of the GC excess at and around 2 GeV, where the GC excess peaks (see figures 3 and 6).
 If past studies have confused part of the halo excess with the extended tail of the GC excess, the true GC excess may be more centrally concentrated than .
@@
 The annihilation cross section estimated from the 20 GeV halo excess ( with GeV for the channel) is marginally consistent with the upper limits obtained by earlier studies on the MW halo region [28, 29, 30, 31, 33].
-The LAT team’s analysis using two years of the LAT data [26] reported without background modeling, for their standard choice of the dark matter density at the solar neighborhood .
+The LAT team’s analysis using two years of the LAT data [28] reported without background modeling, for their standard choice of the dark matter density at the solar neighborhood, .
 It should also be noted that the ROI of the LAT team’s analysis (, ) is much narrower in the Galactic latitude than that of this study.
 The cross section required for the halo excess is larger than the upper limits by dwarf spheroidal galaxies, which are currently considered to be the strongest [12, 13, 14, 15].
-For example, it is several times larger than the 95% containment region of the constraints obtained in ref.
-[13].
+For example, it is several times larger than the 95% containment region of the constraints obtained in a recent study [15].
 However, this tension does not immediately rule out the dark matter interpretation of the 20 GeV halo excess.
 Compared to dwarf galaxies, whose dark matter mass can be more reliably constrained by observations, the density distribution of the MW halo is highly uncertain ([4], see also section 4.3.3 below).
-The dark matter content in dwarf galaxies is also still subject to uncertainties regarding methodology and observed data [65, 66].
-In fact, the dark matter interpretation of the GC GeV excess is also in the same degree of tension with dwarf galaxy constraints, but for the same reason, the dark matter interpretation is not considered to be excluded [8, 13].
+The dark matter content in dwarf galaxies also remains subject to uncertainties regarding methodology and observed data [67, 68].
+In fact, the dark matter interpretation of the GC GeV excess is also in a similar degree of tension with dwarf galaxy constraints, but for the same reason, the dark matter interpretation is not considered to be excluded [8, 15].
 The annihilation cross section determines the era of WIMP decoupling in the early universe, and is therefore related to the relic amount of dark matter in the present universe.
@@
 The uncertainties in the density profile of the MW halo can be described in more detail as follows.
-Since the mass of baryons exceeds that of dark matter in the inner region of the MW ( 10 kpc) [67, 68], we can only extrapolate from the dark matter density at the solar system () assuming a halo density profile, but is uncertain by a factor of a few.
+Since the mass of baryons exceeds that of dark matter in the inner region of the MW ( 10 kpc) [69, 70, 71, 72], we can only extrapolate from the dark matter density at the solar system () assuming a halo density profile.
+In addition to the uncertainty of the density profile, the estimate for has an uncertainty of a factor of about two centered around [4].
+In the case of a smooth density distribution (), a threefold uncertainty in the dark matter density at 10∘–20∘ from the GC changes the annihilation flux by nearly an order of magnitude.
 There are several other analytic functions proposed for the halo density profile, in addition to NFW, and the Einasto profile [73], which is known to fit numerical simulations as well as NFW, is particularly interesting for the halo excess.
 When is fixed, this profile predicts a higher density than NFW at around 10∘ ( kpc) from the GC, and a shallower profile than NFW near the GC (see figure 2.3 of [4]), which are in agreement with the trends suggested by the 20 GeV halo excess.
+Some reports imply that a more core-like density profile than NFW better matches observational data for the MW halo [71, 72].
 In addition to the uncertainty in the fitting formulae for the mean halo density profile in numerical simulations, individual halo profiles also vary depending on their formation and merger history.
-Furthermore, in the central region of a halo, the dark matter density profile may be affected by complicated baryon dynamics, and this effect is not yet fully understood [70, 71].
-In the case of a smooth density distribution (), a threefold uncertainty in the estimate of dark matter density in the 10∘–20∘ region from the GC changes the halo flux by nearly an order of magnitude.
+Furthermore, in the central region of a halo, dark matter density profile may be affected by complicated baryon dynamics, and this effect is not yet fully understood [74, 75].
 Finally, the effect of subhalos or any substructures in the MW halo should not be overlooked.
@@
 Theoretical predictions suggest that the EGB flux by annihilation is comparable to that from the MW halo, with the peak photon energy redshifted to about half [24].
-The observed EGB in the 1–1000 GeV range is thought to contain various astronomical components, such as active galactic nuclei and star-forming galaxies, making it difficult to identify the dark matter component unambiguously.
+The observed EGB in the 1–1000 GeV range is thought to contain various astronomical components, such as active galactic nuclei and star-forming galaxies [25], making it difficult to identify the dark matter component unambiguously.
 A recent study suggests that a significant contribution to the EGB from annihilation with the WIMP parameters of the 20 GeV halo excess is not ruled out [27].
@@
 Annihilation gamma rays from the MW halo and extragalactic halos, whose spectral peak is around 10–20 GeV, may be a candidate for such a component.
-In fact, this was the original motivation behind this study to search for signs of dark matter annihilation.
+In particular, the emission from the MW halo is expected to be relatively stronger at higher photon energies because it is not subject to intergalactic absorption, and some of it may have been confused with the isotropic background radiation.
+In fact, this was the original motivation behind this study to search for signs of dark matter annihilation from the MW halo.
 5 Summary and Discussion
@@
 Looking at the maps of the halo model plus fit residual, the shapes are nearly spherically symmetric in all energy bins where the halo component is detected, indicating that the halo model is a good description of the observed map.
-Various types of systematic uncertainties are examined, including the GALPROP model parameters, the energy bin used to create the Fermi bubble template, and the morphology of the bubble template.
+Various types of systematic uncertainties are examined, including the GALPROP model parameters and the morphology of the Fermi bubble template.
 Due to the strong diffuse emission from cosmic-ray interactions, the systematic uncertainty of the halo component flux becomes larger at lower photon energies.
@@
 Detection of annihilation signals from other objects or regions with consistent WIMP parameters will be crucial for the final confirmation.
-Gamma-ray observations of dwarf galaxies in the MW halo are particularly interesting from this perspective.
-Given the tension with the constraints from dwarf galaxies, a modest improvement in the sensitivity of future observations should eventually lead to the detection of annihilation gamma rays.
+Gamma-ray observations of dwarf galaxies in the MW halo are fascinating from this perspective.
+Given the tension between the MW halo excess and the dwarf galaxy constraints, a modest improvement in the sensitivity of future observations should eventually lead to the detection of annihilation gamma rays.
 It is noteworthy that interesting excesses have already been observed in several dwarf galaxies, even if they have not been definitively detected [15].
 The largest excess has been reported in Reticulum II by several studies (typically –) [77, 78, 79, 80, 15], and the WIMP mass estimated by the latest study [15] is interestingly similar to that for the MW halo excess found by the present study.
-Another important avenue is the search for line gamma-ray emission.
+Another important avenue is to search for line gamma-ray emission.
 Line emissions at the photon energy equal to the WIMP mass (0.3–0.8 TeV for the channel) are expected, although the intensity will depend on the branching ratio of annihilation into two photons (or perhaps ).
 In this energy range, ground-based atmospheric Cherenkov telescopes are more sensitive than the Fermi LAT, and further significant improvements in sensitivity are expected shortly by the Cherenkov Telescope Array Observatory (CTAO) [81].
-Dark matter search using neutrinos is also interesting.
-Though the current neutrino constraints [4, 78, 79, 80] are 1–2 orders of magnitude (depending on the annihilation channels) weaker than the mass and cross section suggested in this work, future improvements in the sensitivity of neutrino detectors may enable independent verification.
+Dark matter searches using neutrinos are also intriguing.
+Though the current neutrino constraints [4, 82, 83, 84] are 1–2 orders of magnitude (depending on the annihilation channels) weaker than the WIMP parameters suggested in this work, future improvements in the sensitivity of neutrino detectors may enable independent verification.
 Since the gamma-ray to neutrino flux ratio can be calculated using the Standard Model of particle physics, the neutrino flux from the Fermi halo can be predicted without uncertainty about the dark matter density profile.
 Acknowledgments
-The author thanks Tsunefumi Mizuno for useful discussion and advice. The author was supported by the JSPS/MEXT KAKENHI Grant Number 18K03692.
+The author thanks Tsunefumi Mizuno for the useful discussion and advice. The author was supported by the JSPS/MEXT KAKENHI Grant Number 18K03692.